# The inducibility of Turán graphs

### v1

@@ 5 Exact results
Let $x\coloneqq a_{k}$ and $y\coloneqq a_{1}$. Then $y-1>x\geq t$. Suppose for a contradiction that both $I(F,G)\geq I(F,G^{\prime})$ and $I(F,G)\geq I(F,G^{\prime\prime})$ hold. Define
$A\coloneqq r\sum_{(i_{1},\dots,i_{r-1})\in([k-1]\setminus\{1\})_{r-1}}\tbinom{a_{i_{1}}}{t}\cdots\tbinom{a_{i_{r-1}}}{t},\quad B\coloneqq r(r-1)\sum_{(i_{1},\dots,i_{r-2})\in([k-1]\setminus\{1\})_{r-2}}\tbinom{a_{i_{1}}}{t}\cdots\tbinom{a_{i_{r-2}}}{t}$.
Since $a_{i}\geq t$ for all $i$ and $k\geq r\geq 2$, the sum defining $B$ contains at least one term (as $k-1\geq r-2$), and each term is positive because $\binom{a_{i_{j}}}{t}\geq 1$; therefore $B>0$. By (4), we obtain the two inequalities
@@
$\alpha_{1}=\tbinom{x}{t}+\tbinom{y}{t}-\tbinom{x+1}{t}-\tbinom{y-1}{t},~\alpha_{2}=\tbinom{x+y}{t}-\tbinom{x}{t}-\tbinom{y}{t},~\beta_{1}=\tbinom{x+1}{t}\tbinom{y-1}{t}-\tbinom{x}{t}\tbinom{y}{t},~\beta_{2}=\tbinom{x}{t}\tbinom{y}{t}$.
Observe that since $y>x\geq t\geq 2$, we have
$\alpha_{1}=\tbinom{y-1}{t-1}-\tbinom{x}{t-1}>0,\quad\alpha_{2}=\sum_{i=1}^{t-1}\tbinom{x}{i}\tbinom{y}{t-i}>0,\quad\beta_{1}=\tfrac{y-x-1}{t}\tbinom{x}{t-1}\tbinom{y-1}{t-1}>0,\quad\beta_{2}>0.$
Combining inequalities (43) and (44) and using $B>0$, we obtain the necessary condition
@@
which, after rearrangement, is equivalent to
@@
Consequently, we obtain
$y\tbinom{y-1}{t}-(x+1)\tbinom{x}{t}\geq(y-x-1)\tbinom{x+y}{t}$.
However, since $z\mapsto\binom{z}{t}$ is strictly convex on $t,\infty)$ for $t\geq 2$, Jensen’s inequality together with the convex combination $\frac{x+1}{y}\cdot x+\frac{y-x-1}{y}\cdot(x+y)=y-1$ gives
$\tfrac{x+1}{y}\tbinom{x}{t}+\tfrac{y-x-1}{y}\tbinom{x+y}{t}>\tbinom{y-1}{t}$,
which is equivalent to $y\tbinom{y-1}{t}-(x+1)\tbinom{x}{t}<(y-x-1)\tbinom{x+y}{t}$, contradicting the previous inequality. This completes the proof of Claim 5.4. ∎
By Claim 5.4, if $a_{1}-a_{k}>1$, then $G$ cannot be extremal. Therefore, in any extremal complete multipartite graph we must have $a_{1}-a_{k}\leq 1$; i.e., all part sizes differ by at most 11. Consequently, every extremal graph is isomorphic to a Turán graph $T_{k}(n)$. This completes the proof of Theorem 5.3. ∎
@@ 6 Concluding remarks
For all $\ell>r\geq 3$, characterize the complete $r$-partite graphs $F$ on $\ell$ vertices for which $\mathrm{OPT}(F)=\{\bm{m}_{r,\ell}\}$.
Assume that $F=K_{a_{1},\ldots,a_{r}}$ is a complete $r$-partite graphs $F$ on $\ell$ vertices. Then a necessary condition for $F$ satisfying $\mathrm{OPT}(F)=\{\bm{m}_{r,\ell}\}$ is
$\binom{\ell}{2}\geq m_{r,\ell}\sum_{k=1}^{r}\binom{a_{k}}{2}$. (45)

### v2

@@ 5 Exact results and uniqueness of the extremal construction
Let $x\coloneqq a_{k}$ and $y\coloneqq a_{1}$, noting that $y-1>x\geq t$. Define
$A$ $\coloneqq\begin{cases}0,&\quad\text{if}\quad r=2,\\[3.99994pt] r\cdot\sum_{(i_{1},\dots,i_{r-1})\in([2,k-1])_{r-1}}\tbinom{a_{i_{1}}}{t}\cdots\tbinom{a_{i_{r-1}}}{t},&\quad\text{if}\quad r\geq 3,\end{cases}$
$B$ $\coloneqq\begin{cases}2,&\quad\text{if}\quad r=2,\\[3.99994pt] r(r-1)\cdot\sum_{(i_{1},\dots,i_{r-2})\in([2,k-1])_{r-2}}\tbinom{a_{i_{1}}}{t}\cdots\tbinom{a_{i_{r-2}}}{t},&\quad\text{if}\quad r\geq 3.\end{cases}$
Since $\min_{i\in[k]}a_{i}\geq t$ and $k\geq r$, it follows directly from the definition that $B>0$.
Suppose for a contradiction that both $I(F,G)\geq I(F,G^{\prime})$ and $I(F,G)\geq I(F,G^{\prime\prime})$ hold. Applying (4) and then simplifying, we obtain
@@
$\alpha_{1}\coloneqq\tbinom{x}{t}+\tbinom{y}{t}-\tbinom{x+1}{t}-\tbinom{y-1}{t},\quad\alpha_{2}\coloneqq\tbinom{x+y}{t}-\tbinom{x}{t}-\tbinom{y}{t},\quad\beta_{1}\coloneqq\tbinom{x+1}{t}\tbinom{y-1}{t}-\tbinom{x}{t}\tbinom{y}{t},\quad\beta_{2}\coloneqq\tbinom{x}{t}\tbinom{y}{t}$.
Since $y>x\geq t\geq 2$, we have
$\alpha_{1}=\tbinom{y-1}{t-1}-\tbinom{x}{t-1}>0,\quad\alpha_{2}=\sum_{i\in[t-1]}\tbinom{x}{i}\tbinom{y}{t-i}>0,\quad\beta_{1}=\tfrac{y-x-1}{t}\tbinom{x}{t-1}\tbinom{y-1}{t-1}>0,\quad\beta_{2}>0.$
Combining this with (39), (40), and $B>0$, we obtain
@@
which after rearranging yields
@@
It follows that
$y\tbinom{y-1}{t}-(x+1)\tbinom{x}{t}-(y-x-1)\tbinom{x+y}{t}\geq 0.$ (41)
Let $\phi\colon[t,\infty)\to\mathbb{R}$ be the function defined by $\phi(z)=\binom{z}{t}$ for every $z\in[t,\infty)$. Since $\phi$ is strictly convex on $t,\infty)$ for $t\geq 2$, it follows from Jensen’s inequality that
$\tfrac{x+1}{y}\tbinom{x}{t}+\tfrac{y-x-1}{y}\tbinom{x+y}{t}>\tbinom{\frac{x+1}{y}\cdot x+\frac{y-x-1}{y}\cdot(x+y)}{t}=\tbinom{y-1}{t}$.
However, this is a contradiction to (41). This completes the proof of Claim 5.7. ∎
It follows from Claim 5.7 that $G$ is not extremal, contradicting the assumption that $I(F,G)=I(F,n)$. This completes the proof of Theorem 5.6. ∎
@@ 6 Concluding remarks
For all integers $\ell>r\geq 3$, characterize the complete $r$-partite graphs $F$ on $\ell$ vertices for which $\mathrm{OPT}(F)=\{\bm{m}_{r,\ell}\}$.
We remark that a necessary condition for a complete $r$-partite graph $F=K_{a_{1},\ldots,a_{r}}$ on $\ell$ vertices satisfying $\mathrm{OPT}(F)=\{\bm{m}_{r,\ell}\}$ is given by the following inequality, which is slightly weaker than that in Lemma 5.3:
$\binom{\ell}{2}\geq m_{r,\ell}\sum_{k\in[r]}\binom{a_{k}}{2}$. (42)